# On the geometry of spaces of filtrations on local rings

### v1

@@ 6.1. Relation to global results
As was mentioned earlier, our definition for the metric $d_{1}$ is inspired by the Darvas metric $d_{1}$ on $\mathcal{H}(X,\omega)$, the space of smooth Kähler potentials of a compact Kähler manifold $(X,\omega)$, introduced in [Dar15], and is a local analogue of the metric $d_{1}$ on $R=R(X,-rK_{X})$, the section ring of a Fano manifold $X$, introduced in [BJ21]. We now explain an approach to equip the global object with a structure in our setting via the cone construction. We work over the complex numbers $\mathbb{C}$ for simplicity.
Let $(V,L)$ be a smooth polarized variety, that is, $V$ is a smooth projective variety and $L$ is an ample line bundle on $V$. Let $R(V,L)\coloneqq\oplus_{m=0}^{\infty}H^{0}(V,mL)=\oplus_{m=0}^{\infty}R_{m}$, then we know that $R_{0}=\mathbb{C}$. We may replace $L$ by $rL$ for some $r\in\mathbb{Z}_{>0}$ such that $R$ is generated by $R_{1}$ as an $R_{0}$-algebra, and in the sequel, we will always assume $L$ is positive enough such that the above condition is satisfied. In this case, the cone $C(V,L)\coloneqq\mathrm{Spec}R(V,L)$ has a normal, isolated singularity at the vertex 00 defined by $R_{+}(V,L)\coloneqq\oplus_{m>0}R_{m}$.
@@
for $t\in[0,+\infty)$. Denote the corresponding norm by $\chi_{t}$, then $\chi_{0}=\mathrm{ord}_{V}$ by definition, and for any $s\in R$, we have $\lim_{t\to+\infty}\chi_{t}(s)=v(s)$. For any $t\in[0,+\infty)$, by definition we have $\mathfrak{m}^{\lceil(1+t)\lambda\rceil}\subset\mathfrak{a}_{\lambda,t}$. Moreover, since $\mathcal{F}_{v}$ is linearly bounded, there exists $C\in\mathbb{R}_{>0}$ such that for any $s\in R_{l}$, we have $v(s)\leq Cl$. So $s\in\mathfrak{a}_{\lambda,t}$ will force $\mathrm{ord}_{V}(s)\geq\frac{(1+t)\lambda}{1+C}$, that is, $\mathfrak{a}_{\lambda,t}\subset\mathfrak{m}^{\lfloor\lambda(1+t)/(1+C)\rfloor}$. This shows that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}$.  [^5] Indeed $\mathfrak{a}_{\bullet,t}=\mathfrak{a}_{\bullet}(v_{t})$ for some $v\in\mathrm{Val}^{<+\infty}_{X,x}$.
Thus we may view $v\in\mathrm{Val}^{<+\infty}_{V}$ as an element sitting on the boundary of $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$. We do not go further in this direction. See also [Fin23] and the references therein for some related results.
@@ 6.2. Examples
In this section we consider several examples. Note that we only work with linearly bounded filtrations, or equivalently, filtrations with positive multiplicity.
@@ 6.2.1. The metric space $(\mathrm{Fil}^{s},d_{1})$ in low dimensions
If $R$ is a DVR, then $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\cong\mathbb{R}_{>0}$. Let $\pi\in\mathfrak{m}$ a uniformizer. For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, let $c\coloneqq\mathrm{ord}_{\pi}(\mathfrak{a}_{\bullet})\in\mathbb{R}_{>0}$. Since $\mathrm{ord}_{\pi}$ is the only divisorial valuation up to scaling, for $\lambda\in\mathbb{R}_{>0}$ we have
@@
Similar to Theorem 1.1, one can also define a pseudo-metric $d_{1}$ on the set $\mathcal{I}_{\mathfrak{m}}$ by
@@
By Rees’ Theorem, the Hausdorff quotient of $(\mathcal{I},d_{1})$ can be identified with $(\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}},d_{1})$. Note that with this distance, the inclusion $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\hookrightarrow\mathrm{Fil}^{s}$ is not an embedding but just a contraction map, since in general we only have $\mathrm{e}(\mathfrak{a}\cap\mathfrak{b})\geq\mathrm{e}(\widetilde{\mathfrak{a}^{\bullet}}\cap\widetilde{\mathfrak{b}^{\bullet}})$.
@@
where $A=\mathbb{Q}_{\geq 0}$ or $\mathbb{R}_{\geq 0}$. Clearly the following metric makes sense
@@ 6.3.1. Completeness under the Darvas metric
The answer to the question is likely no, but we do not have a counterexample.
@@ 6.3.2. More structural questions
Question 6.4.
Is $\mathrm{Fil}^{s}$ always contractible? If not, is there a criterion in terms of the algebraic properties of $R$?

### v2

@@ 6.1. Relation to global results
As was mentioned earlier, our definition for the metric $d_{1}$ is inspired by the Darvas metric $d_{1}$ on $\mathcal{H}(X,\omega)$, the space of smooth Kähler potentials of a compact Kähler manifold $(X,\omega)$, introduced in [Dar15], and is a local analogue of the metric $d_{1}$ on $R=R(X,-rK_{X})$, the section ring of a Fano manifold $X$, introduced in [BJ25]. We now explain an approach to equip the global object with a structure in our setting via the cone construction. We work over the complex numbers $\mathbb{C}$ for simplicity.
Let $(V,L)$ be a smooth polarized variety, that is, $V$ is a smooth projective variety and $L$ is an ample line bundle on $V$. Let $R(V,L)\coloneqq\oplus_{m=0}^{\infty}H^{0}(V,mL)=\oplus_{m=0}^{\infty}R_{m}$, then we know that $R_{0}=\mathbb{C}$. We may replace $L$ by $rL$ for some $r\in\mathbb{Z}_{>0}$ such that $R$ is generated by $R_{1}$ as an $R_{0}$-algebra, and in the sequel, we will always assume $L$ is positive enough such that the above condition is satisfied. In this case, the cone $C(V,L)\coloneqq\mathrm{Spec}R(V,L)$ has a normal, isolated singularity at the vertex 0 defined by $R_{+}(V,L)\coloneqq\oplus_{m>0}R_{m}$.
@@
for $t\in[0,+\infty)$. Denote the corresponding norm by $\chi_{t}$, then $\chi_{0}=\mathrm{ord}_{V}$ by definition, and for any $s\in R$, we have $\lim_{t\to+\infty}\chi_{t}(s)=v(s)$. For any $t\in[0,+\infty)$, by definition, we have $\mathfrak{m}^{\lceil(1+t)\lambda\rceil}\subset\mathfrak{a}_{\lambda,t}$. Moreover, since $\mathcal{F}_{v}$ is linearly bounded, there exists $C\in\mathbb{R}_{>0}$ such that for any $s\in R_{l}$, we have $v(s)\leq Cl$. So $s\in\mathfrak{a}_{\lambda,t}$ will force $\mathrm{ord}_{V}(s)\geq\frac{(1+t)\lambda}{1+C}$, that is, $\mathfrak{a}_{\lambda,t}\subset\mathfrak{m}^{\lfloor\lambda(1+t)/(1+C)\rfloor}$. This shows that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}$.  [^4] Indeed $\mathfrak{a}_{\bullet,t}=\mathfrak{a}_{\bullet}(v_{t})$ for some $v\in\mathrm{Val}^{<+\infty}_{X,x}$.
Thus we may view $v\in\mathrm{Val}^{<+\infty}_{V}$ as an element sitting on the boundary of $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$. We do not go further in this direction. See also [Fin25a] and the references therein for some related results.
@@ 6.2. Examples
In this section, we consider several examples. Note that we only work with linearly bounded filtrations, or equivalently, filtrations with positive multiplicity.
@@ 6.2.1. The metric space $(\mathrm{Fil}^{s},d_{1})$ in low dimensions
If $R$ is a DVR, then $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\cong\mathbb{R}_{>0}$. Let $\pi\in\mathfrak{m}$ be a uniformizer. For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, let $c\coloneqq\mathrm{ord}_{\pi}(\mathfrak{a}_{\bullet})\in\mathbb{R}_{>0}$. Since $\mathrm{ord}_{\pi}$ is the only divisorial valuation up to scaling, for $\lambda\in\mathbb{R}_{>0}$ we have
@@
Similar to Theorem 1.1, one can also define a pseudometric $d_{1}$ on the set $\mathcal{I}_{\mathfrak{m}}$ by
@@
[…]
@@
where $A=\mathbb{Q}_{\geq 0}$ or $\mathbb{R}_{\geq 0}$. Clearly, the following metric makes sense.
@@
6.3. The structure of finitely generated filtrations
Recall that a filtration $\mathfrak{a}_{\bullet}$ is finitely generated, or Noetherian, if the associated graded ring $\mathrm{gr}_{\mathfrak{a}_{\bullet}}R\coloneqq\oplus_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is a finitely generated $\kappa$-algebra.
The finite generation of valuation ideals has been extensively studied since [BCHM10]. See [LXZ22, XZ25] for recent progress based on the theory of special complements.
As another application of Proposition 3.15, we show that finitely generated filtrations are dense in $\mathrm{Fil}_{R,\mathfrak{m}}$, which is the local analogue of [BJ25, Corollary 3.19].
Proposition 6.3.
The subset of finitely generated filtrations is dense in $\mathrm{Fil}_{R,\mathfrak{m}}$ with respect to $d_{1}$.
Proof.
The following construction is a local analogue of [Xu25, Definition-Lemma 3.56]. For $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$ and $k\in\mathbb{Z}_{>0}$, let
$\mathfrak{a}_{\lambda,k}\coloneqq\left\{\begin{aligned} &\mathfrak{a}_{\lambda},&\text{ if }\lambda\leq k,\\ &\sum_{\vec{\mu}}\mathfrak{a}_{\mu_{1}}\cdot\ldots\cdot\mathfrak{a}_{\mu_{s}},&\text{ otherwise,}\end{aligned}\right$.
where the sum is taken over all vectors $\vec{\mu}=(\mu_{1},\ldots,\mu_{s})\in\mathbb{R}_{>0}^{s}$ satisfying $\sum_{i}\mu_{i}\geq\lambda$ and $\mu_{i}\leq k$ for any $1,\ldots,s$. Note that the sum is finite since the set of jumping numbers of $\mathfrak{a}_{\bullet}$ contained in $(0,k]$ is finite.
It is easy to check that $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}_{R,\mathfrak{m}}$. Moreover, $\mathrm{gr}_{\mathfrak{a}_{\bullet,k}}R$ is generated by elements of degree at most $k$, and hence is finitely generated again by the finiteness of jumping numbers.
Clearly $\{\mathfrak{a}_{\bullet,k}\}$ is an increasing sequence and $\cup_{k}\mathfrak{a}_{\bullet,k}=\mathfrak{a}_{\bullet}$. So by Proposition 3.15, we know that $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})\to 0$, which finishes the proof. ∎
Remark 6.4.
In general, the saturation of a finitely generated filtration is not finitely generated. Hence, the above proposition says nothing about the structure or distribution of finitely generated saturated filtrations, which remains rather mysterious.
Indeed, in [LX24], Y. Liu and Xu proposed several conjectures regarding the structure of Kollár valuations, which are special classes of finitely generated filtrations. See [ABB+23, LXZ22, Pen25] for some special cases.
Moreover, in view of [LXZ22, Section 4] and [XZ25, Section 4.3], it seems that the specialty of quasi-monomial valuations inside a single dual complex is related to the geodesics.
@@ 6.4.1. Completeness under the Darvas metric
We remark that convergence along decreasing sequences or nets has been studied in the complex analytic setting [Dar17, Proposition 6.11] and the global non-archimedean setting [BJ22, Theorem 12.5]. The answer to the above question is likely no, but we do not have a counterexample.
@@ 6.4.2. More structural questions
Question 6.6.
Is $\mathrm{Fil}^{s}$ always contractible? If not, is the contractibility related to the algebraic properties of $R$?
Statement for conflict of interest. The author states that there is no conflict of interest.